OpenStudy (anonymous):

solve system of equation: x+y=-1 y-z=4 2x+y+5z=-38 I got 13,-14, and 10 is this correct? or is it -10?

OpenStudy (anonymous):

you can always verify your ans by putting the answers in the equations and seeing if the RHS is same as LHS

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay it is a positive 10. Thanks

OpenStudy (anonymous):

Substitute x=-1-y and z=y-4 into 2x+y+5z=-38

OpenStudy (anonymous):

oh okay i will diana thanks :)

OpenStudy (anonymous):

okay now i'm gettin a negative lol

OpenStudy (anonymous):

thanks diana1

OpenStudy (anonymous):

Did you get the answer? y=-4 Find x and z

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good job! :)